# Permutation-Invariant Spectral Learning via Dyson Diffusion

### v1

@@ 4.1 Methodology
We compare to the GNN-based models EDP-GNN (Niu et al., 2020) and GDSS (Jo et al., 2022), as well as graph-transformer-based ConGress and DiGress (Vignac et al., 2022), where among all the models only DiGress uses a data-augmentation trick: it adds certain graph features, including cycle counts and the first $6$ eigenvalues (for details, see Section 5), but this trick is supposedly inessential for building a good model (Vignac et al., 2022). The trick, however, improves the learning of certain features, but not necessarily other subgraph structures (Wang et al., 2025). We thus compare to both the Digress model “without the trick”, i.e., just a graph (Markov chain) diffusion model, and with the data-augmentation trick.
As we evaluated the full spectrum and the published experimental results only showed partial or no information about the spectrum, we need explicit access to the samples. Since not all models (Vignac et al., 2022; Niu et al., 2020) report snapshots, we had to retrain those also on the standard datasets. We report our code and all the samples of all models in Github  [^5] See https://github.com/schwarzTass/DyDM-code.
WL-Bimodal. Here we train the simple bimodal distribution in Fig. 2, consisting of $80\%$ graph A and $20\%$ graph B, which are WL equivalent. The dataset has $N=5^{\prime}000$ random permutations of A and B, and we follow the standard test/train split procedure (Jo et al., 2022; You et al., 2018; Niu et al., 2020) using $80\%$ of the data as train data and the remaining $20\%$ as test data. We performed hyperparameter tuning of the comparison models as described in Appendix N.
Community. Being a standard benchmark (Niu et al., 2020; Jo et al., 2022; You et al., 2018), we include it for comparison. However, due to heavy undersampling, we only test for memorization. From our perspective, memorization is the best that can be tested with said benchmark, and we elaborate on this in Section N.1
Brain. From the human connectome graph we drew $15^{\prime}000$ ego-graphs (i.e. the induced subgraph of neighborhoods) of size $5$ to $10$ vertices Amunts et al. (2013); Rossi & Ahmed (2015). We give details in Appendix M.
@@ 4.2 Results
In Table 1 we report mean distances and marginal Wasserstein distances of the spectra. Explicitly, if $\nu_{\mathrm{test}}$ is the distribution of the spectrum of the test dataset and $\nu_{\mathrm{samp}}$ the distribution of our samples, then the distance between the means in $\mathbb{R}^{n}$ is $\mu(\nu_{\mathrm{samp}},\nu_{\mathrm{test}})=\norm{\mathbb{E}_{\lambda\sim\nu_{\mathrm{samp}}}[\lambda]-\mathbb{E}_{\lambda\sim\nu_{\mathrm{test}}}[\lambda]}_{2}.$ For statistical feasibility, instead of calculating the full Wasserstein distance we use the averaged marginal Wasserstein distance given by $\mathcal{W}_{\mathrm{marg}}(\nu_{\mathrm{samp}},\nu_{\mathrm{test}})=\frac{1}{n}\sum_{k=1}^{n}\mathcal{W}((\nu_{\mathrm{samp}})_{k},(\nu_{\mathrm{test}})_{k}),$ for $(\nu_{\mathrm{samp}})_{k}$ the marginal distribution in dimension $k$ and $\mathcal{W}$ the Wasserstein distance between two one-dimensional distributions. Using these metrics we can evaluate both marginal and high-dimensional effects.
These metrics also reveal the limitations of GNN- and Graph-Transformer-based models on the simple WL-Bimodal dataset as described in Figure 2, we also see that this extends to the real-world benchmark dataset Community Small. DyDM, on the other hand, consistently overcomes these issues. Even if we compare to the model with ad hoc feature augmentation (Digress with “trick”, on which we elaborate in Section 5), DyDM – which does not employ feature augmentation – either improves on or is on par with the feature augmented DiGress. We demonstrate that this performance still holds when working on large datasets, such as the $15^{\prime}000$ ego-graphs from the Brain dataset.
@@ 5 Related Work
The spectrum carries key features of graphs, and as such has been considered before, however, only as either auxiliary features for data augmentation (Vignac et al., 2022), or in a way which makes any remaining $\Theta(n^{2})$ degrees of freedom inaccessible, as we outline below.
@@
Data augmentation. Vignac et al. (2022) acknowledge the importance of spectra of graphs and add plenty of auxiliary features, among them the first $6$ eigenvalues of the Graph Laplacian as graph-level features and the first $2$ non-zero eigenvectors as vertex-level features. This engineering trick is indeed helpful, yet it is a trick; as we show in Table 1 these auxiliary features are necessary for DiGress to provide a good model. This is in stark contrast to DyDM, where we build on an analytical expression of the evolution of all eigenvalues during diffusion of a symmetric matrix.
@@
Spectral OU method. The use of spectral information for a graph diffusion model has been explored in recent work by Luo et al. (2024), which, however, performs an OU diffusion solely on the spectrum as if the eigenvectors were not impacted by diffusion, hence making any recovery of remaining information in the form of the eigenvectors impossible. The authors argue that the spectrum contains much information, so that upon sampling the spectrum from the OU-based model, the eigenvectors are simply sampled by taking the eigenvectors of a training sample chosen uniformly at random. However, this strategy is very limited, in that it allows only for a diffusion in $n$ parameters, losing $\Theta(n^{2})$ degrees of freedom irrecoverably. Since we learn in DyDM the spectrum of an OU diffusion on the entire graph, the remaining information remains accessible, see Theorem 3.2. This allows learning the eigenvectors beyond sampling uniformly from training data (Luo et al., 2024).
6 Extensions
A key property of DyDM is that it learns the spectrum while retaining all remaining $\Theta(n^{2})$ degrees of freedom accessible through Eigenvector-SDE. Since Dyson-BM decouples from Eigenvector-SDE, future work could implement a model of Eigenvector-SDE and generate entire graphs. To parameterize Dyson-BM, a graph-transformer model such as Jo et al. (2022) could be used; since Eigenvector-SDE is conditioned on the eigenvalue path which we learn without WL-blindness, we assume that this biases paths sufficiently far from each other to mitigate any WL-problems in classical graph-transformer based methods. Moreover, instead of learning the spectrum of the adjacency matrix $\lambda(A(G))$, one could also learn the combinatorial graph Laplacian $\lambda(L(G))$. Finally, we considered graphs with real (or integer) valued weights, while Dyson-BM is well-defined in other algebras. Hence, generalizations to complex-weighted graphs (Tian & Lambiotte, 2024) are straightforward: Instead of working on $\mathrm{Sym}(\mathbb{R}^{n\times n})$, one works on Hermitian ensembles. Similarly, matrices over the algebra of real quaternions may be considered (Dyson, 1962).
@@ 6 Extensions
Beyond graphs. Since the Dyson SDE is defined on the domain of $\mathrm{Sym}(\mathbb{R}^{n\times n})$ (and even Hermitian matrices), it could be applied beyond graphs to other data. For instance, if correlations between $n$ points is measured in form of covariance matrices, DyDM could learn the spectrum and thereby quantify how strong the data clusters into (low-dimensional) principal components (Chen et al., 2015; Estavoyer & François, 2022; Hess, 2000).
@@ 7 Conclusion
Leveraging the analytical insights offered by Dyson’s Brownian Motion, we introduced DyDM, a diffusion model for spectral learning. Using techniques from Random Matrix Theory, we derived an evolution equation for the corresponding eigenvectors, which renders the remaining information about the underlying matrix available. On the domain of graphs, we demonstrated the struggle of existing learning architectures, e.g., GNN- and graph-transformer-based models. Building on the analytical insights, we decompose the dynamics, such that the spectral part is not constrained by inductive bias, thereby expanding the scope of suitable learning architectures. This way, DyDM can learn the spectrum (even of challenging graph families) without constraining to permutation-equivariant networks. This eliminates the hallucination, so that DyDM learns the distributions without struggle and without any need for data augmentation, as demonstrated experimentally. We hope that this approach opens a new direction in enforcing the inductive bias beyond the learning architecture in graph diffusion models.
@@ Acknowledgments
We thank Jakiw Pidstrigach and Marc Roth for helpful discussions. Financial support from the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation program (Grant Agreement No. 101086182 to A. G.), the Rhodes Trust (to T. S.) and the Swiss National Science Foundation (to C.K. grant number 235409) are acknowledged. This work used the Scientific Compute Cluster at GWDG, the joint data center of Max Planck Society for the Advancement of Science (MPG) and University of Göttingen. In part funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 405797229.
@@
The Appendix is structured as follows. We first derive the tractable loss $\tilde{L}(\theta)$ in Appendix A. We then provide proofs of the Equation Dyson-BM in Appendix B as well as for the Eigenvector-SDE in Appendix C. We then explain the applicability of Anderson’s time reversal (Appendix D), and provide a time-rescaling of Dyson-BM in Appendix E. In Appendix F, we provide the adaptive step size controller, followed by an explanation of the shooting mechanism (Appendix G), sampling procedure (Appendix H), and engineering details (Appendix I). We explain the challenges of Dyson’s Brownian Motion which we overcame with DyDM in Appendix J. We present our theoretical argument for using inductive bias in Appendix K and prove the Lemma on WL equivalence of $k$-regular graphs in Appendix L. For the experiments, we first elaborate on the datasets (Appendix M) followed by an explanation of our extensive benchmarking in Appendix N, including a comment about undersampling in some benchmark datasets in Section N.1. We show the learning dynamics of $4$ different runs of EDP-GNN in Appendix O.
@@ Appendix A Making the loss tractable
In this section, we deduce a general loss formula for an SDE on $\mathbb{R}^{d}$ that will be applied to (Dyson-BM). We consider the process $X=(X(t))_{t\geq 0}$ determined by the SDE
@@
The canonical loss arising from $s(y,t)=\gradient_{M(t)}\log p_{t}(M)$ is
$L^{\prime}(\theta)=\operatorname{\mathbb{E}}_{t\sim\mathrm{Unif}[0,T],M(0)\sim p_{0},M(t)\sim p_{t\mid 0}(\cdot|M(0))}\left[\norm{s_{\theta}(M(t),t)-\gradient_{M(t)}\log p_{t}(M(t))}^{2}_{2}\right].$ (8)
It is a well-known fact that $L^{\prime}(\theta)=L(\theta)+\mathrm{const}(\theta)$. We will first explain how the loss $L^{\prime}(\theta)$ from (8) and $\tilde{L}(\theta)$ from (6) are the same up to a constant, that is $L^{\prime}(\theta)=\tilde{L}(\theta)+\mathrm{const}(\theta)$, which therefore results in the same gradient descent as with $L(\theta)$.
We generalize the loss from (8) to weighing the time by a function $\eta$. So we consider the following generalized loss
$L(\theta)\coloneqq$ $\frac{1}{T}\int_{0}^{T}\eta(t)\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\norm{s_{\theta}(y,t)-\gradient_{y}\log(p_{t}(y))}_{2}^{2}\;p_{t,0}(y,x)\,\mathrm{d}y\mathrm{d}x\mathrm{d}t$
$=$ $\frac{1}{T}\int_{0}^{T}\eta(t)\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\norm{s_{\theta}(y,t)-\gradient_{y}\log(p_{t}(y))}_{2}^{2}\;p_{t\mid 0}(y\mid x)\mathrm{d}yp_{0}(x)\mathrm{d}x\mathrm{d}t,$ (9)
where $\eta(t)$ is some weighting function with $\int_{0}^{T}\eta(t)=1$. We wrote it in the above form with $p_{t\mid 0}$ denoting the conditional density, since the three integrals can be replaced by $\operatorname{\mathbb{E}}_{t\sim U{[0,T]}}\left[\cdots\operatorname{\mathbb{E}}_{x\sim p_{0}}\left[\operatorname{\mathbb{E}}_{y\sim p_{t\mid 0}(\cdot\mid x)}\left[\norm{\cdots}_{2}^{2}\right]\right]\right]$, which we can sample from if we assume (1) sample access to $p_{0}$, (2) known density at any $t$ given a dirac-delta $p_{0}$, (3) the term in the norm is tractable. (1) is assumed by the problem definition, (2) is known for an Ornstein-Uhlenbeck forward SDE, (3) can be solved by realizing that there is an equivalent loss function $\hat{L}(\theta)=L(\theta)+\rm const(\theta)$ where $p_{t}$ inside the norm is converted to a $p_{t\mid 0}$, which is known (gaussian density) in an OU setting.
@@
$\norm{s_{\theta}(y,t)-\gradient_{y}\log(p_{t}(y))}_{2}^{2}=\norm{s_{\theta}(y,t)}_{2}^{2}+\norm{\gradient_{y}\log(p_{t}(y))}_{2}^{2}-2s_{\theta}(y,t)^{T}\gradient_{y}\log(p_{t}(y))$
We will rewrite the mixed term $s_{\theta}(y,t)^{T}\gradient_{y}\log(p_{t}(y))$ of the $L^{2}$ norm in such a way that the first quadratic term $\norm{s_{\theta}(y,t)}_{2}^{2}$ remains unchanged and the second quadratic term $\norm{\gradient_{y}\log(p_{t}(y))}_{2}^{2}$ is constant in $\theta$. First, we rewrite
$\frac{1}{T}\int_{0}^{T}\eta(t)\int_{\mathbb{R}^{n}\times\mathbb{R}^{n}}s_{\theta}(y,t)^{T}\gradient_{y}\log(p_{t}(y))p_{t,0}(y,x)\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}t$ (10)
$=$ $\frac{1}{T}\int_{0}^{T}\eta(t)\int_{\mathbb{R}^{n}}s_{\theta}(y,t)^{T}\gradient_{y}\log(p_{t}(y))p_{t}(y)\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}t$
@@
$=\int_{\mathbb{R}^{n}}\gradient_{y}\log p_{s|s^{\prime}}(y|z)p_{s,s^{\prime}}(y,z)\mathop{}\!\mathrm{d}z.$
@@
$\tilde{L}(\theta)=\frac{1}{T}\int_{0}^{T}\eta(t+h)\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\norm{s_{\theta}(y,t+h)-\gradient_{y}\log(p_{t+h|t}(y|z))}_{2}^{2}p_{t+h|t}(y|z)\mathop{}\!\mathrm{d}y\,p_{t}(z)\mathop{}\!\mathrm{d}z\mathop{}\!\mathrm{d}t$ (13)
@@
$\tilde{L}(\theta)\approx$ $\frac{1}{N}\sum_{r=1}^{N}\sum_{i=1}^{k^{(r)}}\frac{t_{i}^{(r)}-t_{i-1}^{(r)}}{T}\eta(t_{i}^{(r)})\norm{s_{\theta}(x_{i}^{(r)},t_{i}^{(r)})-\gradient_{y}\left[\log(p_{t_{i}^{(r)}\mid t_{i-1}^{(r)}}(x_{i}^{(r)}\mid x_{i-1}^{(r)}))\right]}_{2}^{2}$
@@
We finally approximate the incremental score function $\gradient_{x_{i}^{(r)}}\left[\log(p_{t_{i}^{(r)}\mid t_{i-1}^{(r)}}(x_{i}^{(r)}\mid x_{i-1}^{(r)}))\right]$ as follows. If $h>0$ is a small time step, we can approximate the conditional random variable $X_{t+h}\mid\mathcal{F}_{t}$ with $x:=X_{t}$ by
@@
$\gradient_{y}\log\left(p_{t+h|t}(y|x)\right)$ $\approx-\frac{y-x-a(t,x)h}{b^{2}(t,x)h}.$ (17)

### v2

@@ 4.1 Methodology
We compare to the GNN-based models EDP-GNN [8] and GDSS [9], as well as graph-transformer-based ConGress and DiGress [10], where among all the models only DiGress uses a data-augmentation trick: it adds certain graph features, including cycle counts and the first $6$ eigenvalues (for details, see Section˜5), but this trick is supposedly inessential for building a good model [10]. The trick, however, improves the learning of certain features, but not necessarily other subgraph structures [29]. We thus compare to both the DiGress model “without the trick”, i.e., just a graph (Markov chain) diffusion model, and with the data-augmentation trick. As we evaluated the full spectrum while the published experimental results only showed partial or no information about the spectrum, we need explicit access to the samples. Since not all models [10, 8] report snapshots, we had to retrain those also on the standard datasets. We test DyDM on the following datasets, taking recent advice on limitations of certain graph benchmarks into account [30].
WL-Bimodal. We train the simple bimodal distribution in Fig. 2, consisting of $80\%$ graph A and $20\%$ graph B, which are WL equivalent. The dataset has $N=5,000$ random permutations of A and B, and we follow the standard test/train split procedure [9, 31, 8] using $80\%$ of the data as train data and the remaining $20\%$ as test data. For hyperparameter tuning of the comparison models, see App. Q.
Community. Being a standard benchmark [31, 8, 9], we include it for comparison. However, due to heavy undersampling, we can only test for memorization. From our perspective, memorization is the best that can be tested with said benchmark, and we elaborate on this in App. Q.1.
Brain. From the human connectome graph we drew $15,000$ ego-graphs (i.e. the induced subgraph of neighborhoods) of size $5$ to $10$ vertices [32, 33]. Crucially, this dataset demonstrates scalability to a number of graphs large enough for high statistical fidelity (faithful representation of the underlying distribution) in test and train set. For details, see App. P.
@@ 4.2 Results
In Table˜1 we report mean distances and marginal Wasserstein distances of the spectra. Explicitly, if $\nu_{\mathrm{test}}$ is the distribution of the spectrum of the test dataset and $\nu_{\mathrm{samp}}$ the distribution of our samples, then the distance between the means in $\mathbb{R}^{n}$ is $\mu(\nu_{\mathrm{samp}},\nu_{\mathrm{test}})\!=\!\norm{\mathbb{E}_{\lambda\sim\nu_{\mathrm{samp}}}[\lambda]\!-\!\mathbb{E}_{\lambda\sim\nu_{\mathrm{test}}}[\lambda]}_{2}$}. For statistical feasibility, instead of calculating the full Wasserstein distance we use the averaged marginal Wasserstein distance given by $\mathcal{W}_{\mathrm{marg}}(\nu_{\mathrm{samp}},\nu_{\mathrm{test}})=\frac{1}{n}\sum_{k=1}^{n}\mathcal{W}((\nu_{\mathrm{samp}})_{k},(\nu_{\mathrm{test}})_{k}),$ for $(\nu_{\mathrm{samp}})_{k}$ the marginal distribution in dimension $k$ and $\mathcal{W}$ the Wasserstein distance between two 1d distributions. With these metrics we can evaluate both marginal and high-dimensional effects.
These metrics also reveal the limitations of GNN- and Graph-Transformer-based models on the simple WL-Bimodal dataset, as described in Figure˜2. We also see that this extends to the real-world benchmark dataset Community Small. DyDM, on the other hand, consistently overcomes these issues. Even if we compare to the model with ad hoc feature augmentation (DiGress with “trick”, on which we elaborate in Section˜5), DyDM – which does not employ feature augmentation – either improves on or is on par with the feature augmented DiGress. We demonstrate that this performance still holds when working on large datasets, such as the $15,000$ ego-graphs from the Brain dataset.
With the learned eigenvectors in App. S, we demonstrate in Table 5 that the remaining information remains indeed accessible and viable to generate accurate and valid eigenvectors. More precisely, our method performs best in all metrics except the trivial degree distribution in the WL-case – which aligns with our discussion of the degree distribution being uniform on the WL-equivalence class.
@@ 5 Related Work
The spectrum carries key features of graphs [34], so that spectral methods have generally proven fruitful in graph research, exploiting information encoded in the spectrum, e.g. for graph comparison [35], or in dominant eigenvectors, such as in clustering, community detection [36, 37], and network embedding [38]. For graph diffusion models, however, the spectrum has been used only as either auxiliary features for data augmentation [10], or in a way which makes any remaining $\Theta(n^{2})$ degrees of freedom inaccessible, as we outline below.
@@
Data augmentation. Ref. [10] acknowledges the importance of spectra of graphs and adds several auxiliary features, among them the first $6$ eigenvalues of the Graph Laplacian as graph-level features and the first $2$ non-zero eigenvectors as vertex-level features. While the “trick” of adding features is not principled, we show in Table˜1 that these auxiliary features are necessary for DiGress to provide a good model. These features are not needed by DyDM, which builds on an analytical expression of the evolution of all eigenvalues during diffusion of a symmetric matrix.
@@
Spectral methods. Spectral methods for graph learning have been employed in a variety of contexts, for instance with GANs [39], which have, however, been outperformed by diffusion-based models [10]. Spectral information for graph diffusion models has been explored in recent work, however, none of them considers how a diffusion of the spectrum impacts the remaining information (eigenvectors), leading to either a model that cannot guarantee orthogonality of the eigenvectors [40], or a model solely on the spectrum as if the eigenvectors were not impacted by diffusion [17], hence making any recovery of remaining information in the form of the eigenvectors impossible. Ref. [17] argues that the spectrum contains much information, so that upon sampling the spectrum from the OU-based model, the eigenvectors are simply sampled by taking the eigenvectors of a training sample chosen uniformly at random. However, this strategy is very limited, as it allows only for a diffusion in $n$ parameters, losing $\Theta(n^{2})$ degrees of freedom irrecoverably. Since we learn in DyDM the spectrum of an OU diffusion on the entire graph, the remaining information remains accessible, see Theorem˜3. This allows learning the eigenvectors beyond sampling uniformly from training data [17]. Our Lie group formulation enforces orthogonality intrinsically, without unnecessarily overparameterizing the degrees of freedom (using n(n1)/2+nn(n-1)/2+n instead of $n(n+1)$) and without the need for rejection sampling [40].
6 Limitations and Extensions
@@
Beyond graphs. Our experiments are limited to graphs. However, since the Dyson SDE is valid for the domain of $\mathrm{Sym}(\mathbb{R}^{n\times n})$, it is applicable beyond graphs. For instance, if correlations between $n$ points are represented as covariance matrices, DyDM could learn the spectrum and thereby quantify the effective dimensionality or concentration of variance along principal components [41, 42, 43].
Complex values. While our equations and numerics are limited to real-valued data, an interesting generalization of Dyson-BM is to the complex domain. Thus, generalizing to complex-weighted graphs [44, 45]: instead of working on $\mathrm{Sym}(\mathbb{R}^{n\times n})$, one works on the space of Hermitian matrices. Similarly, matrices over the algebra of real quaternions may be considered [24].
Emphasis on dominant eigenmodes. By changing the loss function, lower eigenvalue-eigenvector pairs could be penalized stronger, thereby learning the graph’s main modes, determining key structural properties [34].
@@ 7 Conclusion
Leveraging the analytical insights offered by Dyson’s Brownian motion from random matrix theory, we introduced DyDM, a diffusion model for spectral learning. In the graph domain, we demonstrated the limitations of existing architectures, such as GNNs and graph transformers. Building on these insights, we decouple the dynamics so that the spectral component is not constrained by inductive biases, thereby expanding the scope of suitable architectures. This allows DyDM to learn the spectrum – even for challenging graph families – without being restricted to permutation-equivariant networks. This approach eliminates architecture-induced hallucination, enabling DyDM to learn the true distributions accurately and without data augmentation, as demonstrated experimentally. Furthermore, our framework and the eigenvector SDE lay the foundation for several future extensions with applications also beyond graphs. We hope this work opens a new direction for enforcing inductive biases beyond the choice of architecture.
@@ Acknowledgments
We thank Jakiw Pidstrigach and Marc Roth for helpful discussions. Financial support from the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation program (Grant Agreement No. 101086182 to A. G.), the Rhodes Trust and EPSRC Centre for Doctoral Training in Mathematics of Random Systems: Analysis, Modelling and Simulation (EPSRC Grant EP/S023925/1) (to T. S.) and the Swiss National Science Foundation (to C.K. grant number 235409) are acknowledged. This work used the Scientific Compute Cluster at GWDG, the joint data center of Max Planck Society for the Advancement of Science (MPG) and University of Göttingen. In part funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 405797229. We acknowledge the help of LLMs to aid in implementing standard methods.
@@
The Appendix is structured as follows. We first derive the tractable loss $\tilde{L}(\theta)$ in Appendix˜A. We then provide proofs of the (Dyson-BM) SDE in Appendix˜B as well as for the (Eigenvector-SDE) in Appendix˜C. We then explain the applicability of Anderson’s time reversal (Appendix˜D), and provide a time-rescaling of (Dyson-BM) in Appendix˜E. In Appendix˜F, we provide the adaptive step size controller, followed by an explanation of the shooting mechanism (Appendix˜G). We provide a theoretical analysis of the complexity of the numerical update steps for (Dyson-BM) and (Eigenvector-SDE) in dependence of the graph size $n$ in Appendix˜H, followed by a discussion of empirical resource considerations in Appendix˜I. We then explain the sampling procedure (Appendix˜J) and give engineering details (Appendix˜K). We explain the challenges of Dyson’s Brownian motion which we overcame with DyDM in Appendix˜L. We present our theoretical argument for using inductive bias in Appendix˜M and prove the Lemma on WL equivalence of $k$-regular graphs in Appendix˜N. In Appendix˜O, we demonstrate DyDM on the Graph Laplacian, showing its applicability beyond adjacency spectra and show how it can learn properties such as the algebraic connectivity (Fiedler value). For the experiments, we first elaborate on the datasets (Appendix˜P) followed by an explanation of our extensive benchmarking in Appendix˜Q, including a comment about undersampling in some benchmark datasets in Section˜Q.1. We show the learning dynamics of $4$ different runs of EDP-GNN in Appendix˜R. We give mathematical details for the extension to (Eigenvector-SDE) and its time-reversal in Appendix˜S, and demonstrate the recoverability of the remaining information in Table 5 and by plotting the generated graphs in Fig. 12.
@@ Appendix A Making the loss tractable
In this section, we deduce a general loss formula for an SDE on $\mathbb{R}$ that will be applied to (Dyson-BM). We consider the process $X=(X(t))_{t\geq 0}$ determined by the SDE
@@
The canonical loss arising from $s(y,t)=\gradient_{M(t)}\ln p_{t}(M)$ is
$L^{\prime}(\theta)=\operatorname{\mathbb{E}}_{t\sim\mathcal{U}[0,T],M(0)\sim p_{0},M(t)\sim p_{t\mid 0}(\cdot|M(0))}\left[\norm{s_{\theta}(M(t),t)-\gradient_{M(t)}\ln p_{t}(M(t))}^{2}_{2}\right].$ (7)
Note that the difference between $L^{\prime}(\theta)$ and $L(\theta)$ is that we use the gradient of the conditional density $\gradient_{M(t)}\ln p_{t\mid 0}(M(t)|M(0))$ in $L(\theta)$. It is a well-known fact that $L^{\prime}(\theta)=L(\theta)+\mathrm{const}(\theta)$. We will first explain how the loss $L^{\prime}(\theta)$ from (7) and $\tilde{L}(\theta)$ from (5) are the same up to a constant, that is $L^{\prime}(\theta)=\tilde{L}(\theta)+\mathrm{const}(\theta)$, which therefore results in the same gradient descent as with $L(\theta)$.
We generalize the loss from (7) to weighting time by a function $\eta$. So we consider the following generalized loss
$L^{\prime}(\theta)\coloneqq$ $\frac{1}{T}\int_{0}^{T}\eta(t)\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\norm{s_{\theta}(y,t)-\gradient_{y}\ln(p_{t}(y))}_{2}^{2}\;p_{t,0}(y,x)\,\mathrm{d}y\mathrm{d}x\mathrm{d}t$
$=$ $\frac{1}{T}\int_{0}^{T}\eta(t)\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\norm{s_{\theta}(y,t)-\gradient_{y}\ln(p_{t}(y))}_{2}^{2}\;p_{t\mid 0}(y\mid x)\mathrm{d}yp_{0}(x)\mathrm{d}x\mathrm{d}t,$ (8)
where $\eta(t)$ is some weighting function with $\int_{0}^{T}\eta(t)\mathrm{d}t=1$. We wrote it in the above form with $p_{t\mid 0}$ denoting the conditional density, since the three integrals can be replaced by $\operatorname{\mathbb{E}}_{t\sim U{[0,T]}}\left[\cdots\operatorname{\mathbb{E}}_{x\sim p_{0}}\left[\operatorname{\mathbb{E}}_{y\sim p_{t\mid 0}(\cdot\mid x)}\left[\norm{\cdots}_{2}^{2}\right]\right]\right]$, which we can sample from if we assume (1) sample access to $p_{0}$, (2) known density at any $t$ given a dirac-delta $p_{0}$, (3) the term in the norm is tractable. (1) is assumed by the problem definition, (2) is known for an Ornstein-Uhlenbeck forward SDE, (3) can be solved by realizing that there is an equivalent loss function $\hat{L}(\theta)=L(\theta)+\rm const(\theta)$ where $p_{t}$ inside the norm is converted to a $p_{t\mid 0}$, which is known (Gaussian density) in an OU setting.
@@
$\norm{s_{\theta}(y,t)-\gradient_{y}\ln(p_{t}(y))}_{2}^{2}=\norm{s_{\theta}(y,t)}_{2}^{2}+\norm{\gradient_{y}\ln(p_{t}(y))}_{2}^{2}-2s_{\theta}(y,t)^{T}\gradient_{y}\ln(p_{t}(y))$
We will rewrite the mixed term $s_{\theta}(y,t)^{T}\gradient_{y}\ln(p_{t}(y))$ of the $L^{2}$ norm in such a way that the first quadratic term $\norm{s_{\theta}(y,t)}_{2}^{2}$ remains unchanged and the second quadratic term $\norm{\gradient_{y}\ln(p_{t}(y))}_{2}^{2}$ is constant in $\theta$. First, we rewrite
$\frac{1}{T}\int_{0}^{T}\eta(t)\int_{\mathbb{R}^{n}\times\mathbb{R}^{n}}s_{\theta}(y,t)^{T}\gradient_{y}\ln(p_{t}(y))p_{t,0}(y,x)\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}t$ (9)
$=$ $\frac{1}{T}\int_{0}^{T}\eta(t)\int_{\mathbb{R}^{n}}s_{\theta}(y,t)^{T}\gradient_{y}\ln(p_{t}(y))p_{t}(y)\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}t$
@@
$=\int_{\mathbb{R}^{n}}\gradient_{y}\ln p_{s|s^{\prime}}(y|z)p_{s,s^{\prime}}(y,z)\mathop{}\!\mathrm{d}z.$
@@
$\tilde{L}(\theta)=\frac{1}{T}\int_{0}^{T}\eta(t+h)\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\norm{s_{\theta}(y,t+h)-\gradient_{y}\ln(p_{t+h|t}(y|z))}_{2}^{2}p_{t+h|t}(y|z)\mathop{}\!\mathrm{d}y\,p_{t}(z)\mathop{}\!\mathrm{d}z\mathop{}\!\mathrm{d}t$ (12)
@@
$\tilde{L}(\theta)\approx$ $\frac{1}{N}\sum_{r=1}^{N}\sum_{i=1}^{k^{(r)}}\frac{t_{i}^{(r)}-t_{i-1}^{(r)}}{T}\eta(t_{i}^{(r)})\norm{s_{\theta}(x_{i}^{(r)},t_{i}^{(r)})-\gradient_{x_{i}^{(r)}}\left[\ln(p_{t_{i}^{(r)}\mid t_{i-1}^{(r)}}(x_{i}^{(r)}\mid x_{i-1}^{(r)}))\right]}_{2}^{2}$
@@
We finally approximate the incremental score function $\gradient_{x_{i}^{(r)}}\left[\ln(p_{t_{i}^{(r)}\mid t_{i-1}^{(r)}}(x_{i}^{(r)}\mid x_{i-1}^{(r)}))\right]$ as follows. If $h>0$ is a small time step, we can approximate the conditional random variable $X_{t+h}\mid\mathcal{F}_{t}$ with $x:=X_{t}$ by
@@
$\gradient_{y}\ln\left(p_{t+h|t}(y|x)\right)$ $\approx-\frac{y-x-a(t,x)h}{b^{2}(t,x)h}.$ (16)